1. Insoluble salts are often by precipitation reactions.

(a) copy and complete the following precipitation reaction equation.

Mg(NO3)2 + K3PO4 =

(b) Derive a net ionic equation of the reaction (a).
(c) Give the name of precipitate (product).
(d) Based on redox type of reactions, define the meaning of the term “spectator ions”
1
Expert's answer
2020-10-23T10:48:57-0400

(a) 3Mg(NO3)2 + 2K3PO4 = Mg3(PO4)2 + 6KNO3

(b) 3Mg2+ + 6NO3- + 6K+ + 2PO43- = Mg3(PO4)2 + 6K+ + 6NO3-

From here, net ionic equation is as following:

3Mg2+ + 2PO43- = Mg3(PO4)2

(c) Trimagnesium phosphate

(d) Spectator ions are ions that do not participate in the reaction and are present in the solution without changing their form. Here, K+ and NO3- are spectator ions.